OpenStudy (anonymous):

How you do formulas in Geo. for 8th graders.

OpenStudy (anonymous):

What do you mean by "How you do formulas"

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for circles

OpenStudy (anonymous):

Area and circumference?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

Area of circle = \[\pi r^2\] Circumference = \[2\pi r\]

OpenStudy (anonymous):

how do you estimate the diameter of a wheel that makes 9 revolutions and travels 50 feet. Using 22/7 for pie

OpenStudy (anonymous):

d = 2r revolutions = circumference

OpenStudy (preetha):

Length = 50 feet. Divide total length travelled by circumference and solve for r

OpenStudy (preetha):

good job abdulshaheer

OpenStudy (anonymous):

What's the answer

OpenStudy (preetha):

You have to put in the numbers and calculate it. We have shown you how to do it, now do it.

OpenStudy (anonymous):

thx but, rude... much

OpenStudy (preetha):

Sorry, did not mean to be rude,but Ireally think you should put in the numbers. Otherwise you will not be able to do that next time. I apologize.

OpenStudy (anonymous):

thx for saying srry

OpenStudy (preetha):

Good luck on Geometry. You'll get the hang of it.